OpenStudy (anonymous):

An equation is shown below: 9x - y = -2 Explain how you will show all of the solutions that satisfy this equation. Determine three different solutions for this equation. Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent.

OpenStudy (anonymous):

I don't understand this very well, can someone explain to me how to do this?

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

you have to put in a number to equal the value! so 9 * x - y = -2 which equals 9 * 3 - 29 = -2

OpenStudy (anonymous):

IS there multiple chooses?

OpenStudy (anonymous):

No... I need to type the work myself

OpenStudy (anonymous):

Exactly, so I just answered your question. If there is a number and a letter then you times it, then there is a subtraction sign so subtract it and you get your answer. so 9 * 3 - 29 = -2 might be a good example on how to do it

OpenStudy (anonymous):

Did I help you?

OpenStudy (anonymous):

Yes, Thanks though, I didn't really understand what you meant until I reread it a 3rd time, lol I need help with the third part though...

OpenStudy (anonymous):

@mario0012

OpenStudy (anonymous):

oh sorry gimme a second

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

Back, I was doing research. I can't understand this. Sorry